Can you help me find the solution to each system by the addition (elimination) method. 4s - 5t = 18 and 6s + 10t = -8
Answer by drk(1908)

TO solve using elimination, we need all variables on the same side and lined up.
(i) 4s - 5t = 18
(ii) 6s + 10t = -8
multiply (i) by 2. We get
(iii) 8s - 10t = 36
(ii) 6s + 10t = -8
add down and solve for s.
(iv) 14s = 28
s = 2.
Find t.
t = -2
(2, -2)